OpenStudy (anonymous):

PLEASE SOLVE WILL GIVE MEDAL Find the derivative of f(x) = -3/x at x=-4

OpenStudy (anonymous):

\[\frac{ d }{ dx } \frac{ 1 }{ x } \implies \frac{ d }{ dx } x^{-1}\] \[\frac{ d }{ dx } x^n \implies nx^{n-1}\]

OpenStudy (anonymous):

I don't understand what that means @iambatman

OpenStudy (anonymous):

No worries, probably the notation :P, we'll use f'(x), do you know the derivative of 1/x?

OpenStudy (anonymous):

\[f(x) = x^n, f'(x) = nx^n\]

OpenStudy (anonymous):

no I don't know

OpenStudy (anonymous):

Do you know what a derivative is?

OpenStudy (anonymous):

no not really

OpenStudy (kainui):

@faariat What's the limit definition of a derivative?

OpenStudy (anonymous):

\[f'(a) = \lim_{h \rightarrow 0} \frac{ f(a+h)-f(a) }{ h }\] this ring a bell?